# Uncertainty-Aware Data-Efficient AI: An Information-Theoretic Perspective

### v1

@@ I Introduction
  • Quantifying epistemic and predictive uncertainty: In the presence of limited training data, epistemic uncertainty fundamentally limits the precision of any AI model. Uncertainty quantification is thus an essential step to ensure the reliable deployment of AI models in domains such as robotics, telecommunications, and healthcare. We start by reviewing Bayesian learning as a formal framework to quantify epistemic uncertainty in the parameter space, as sketched in Fig. 1(a) (with $\theta$ representing the model parameters). We also cover recent “post-Bayesian” advances including generalized Bayesian learning [11] and martingale posteriors [11, 6]. We then review information-theoretic generalization bounds [9]. As illustrated in Fig. 1(b), generalization bounds offer insights into the scaling of the average error caused by limitations in the training data. As it will be shown, information-theoretic generalization bounds provide a theoretical justification for the use of generalized Bayesian learning. However, these bounds do not provide any actionable quantification of uncertainty for individual inputs. As exemplified in Fig. 1(c), this type of information can be instead obtained via prediction sets calibrated through conformal prediction, which offer distribution-free, finite-sample coverage guarantees [12, 13].
  • Leveraging synthetic data: While uncertainty quantification is a required step whenever training data are in limited supply, in practice one may wish to find ways to augment the dataset with additional data. When real data cannot be collected, or is too expensive to obtain, the only remaining option is to leverage synthetic data. This is potentially feasible by leveraging simulators, or digital twins, of real environments [14, 15, 3] or by adopting powerful general-purpose models such as LLMs [16]. As illustrated in Fig. 1(d), synthetic data can be used both to improve training and to enhance the calibration of a pre-trained model. In this context, we review prediction-powered inference, which combines small labeled samples with abundant synthetic predictions for training [17, 18], as well as synthetic-powered predictive inference, which integrates synthetic data for calibration [19, 20].
@@
The definition of the information-theoretic measures used in this article is provided in Table I, while Table II reports a table of acronyms. A summary of information-theoretic relationships surveyed in this paper can be found in Fig. 2.
@@ II Generalized Bayesian Learning
Classical Bayesian inference provides a principled framework for incorporating prior knowledge and quantifying epistemic uncertainty through posterior distributions over parameters affecting the data distribution. The principle can be directly applied to the training of AI models, in which a posterior distribution over the model parameters (denoted as $\theta$) is inferred by combining prior information with a likelihood function [21]. As shown in Fig. 1(a), this leads to a distribution in the model parameter space, whose extent and shape capture the uncertainty arising from the availability of limited data. In practice, the posterior distribution is approximated using sampling, variational methods, or Laplace approximations [22, 6].
@@
ERM Empirical risk minimization
PAC Probably approximately correct
The validity of the posterior distribution hinges on the adherence of prior and likelihood to the ground-truth data distribution. However, prior distributions are often chosen for their analytical tractability, and typical likelihood models fail to capture aspects such as outliers [11, 23]. To obviate these issues, so-called “post-Bayes” methods have been introduced. These move beyond conventional Bayesian inference, while retaining the key goal of quantifying epistemic uncertainty. After reviewing conventional Bayesian learning, this section provides a brief introduction to generalized Bayesian learning and martingale posteriors as two important representative post-Bayes frameworks.
@@ II-A Bayesian Learning
In standard Bayesian inference of a parameter $\theta$, one starts by choosing a prior $p(\theta)$ and likelihood function $p(z|\theta)$, where $z$ represents a data point. In supervised learning settings, the likelihood $p(z|\theta)$ takes the form of the conditional distribution $p(y|x,\theta)$ with input variables $x$ and output variable $y$. Given a training set $\mathcal{D}=\{z_{1},\ldots,z_{n}\}$, the posterior distribution is given by
@@
Unlike conventional frequentist learning, which returns a single model parameter $\theta$, the posterior distribution (1) directly quantifies epistemic uncertainty through the spread of probability mass in the parameter space. Wide posterior distributions indicate high uncertainty about which parameters best explain the data, signaling insufficient training samples. As the training dataset grows in size, i.e., in the limit $n\to\infty$, consistency results show that, under appropriate conditions, the generalized posterior concentrates around the optimal parameter $\theta^{*}$, as illustrated in Fig. 1(a). Specifically, the optimal parameter $\theta^{*}$ provides the best fit for the true data distribution in terms of the Kullback-Leibler (KL) divergence (see Table I) [24].
Trained predictive or generative models can be obtained by drawing one or more samples from the distribution $p(\theta|\mathcal{D})$, typically using Monte Carlo methods [22]. For example, the unadjusted Langevin algorithm produces the sequence of samples
@@
The quantity (5) is a measure of the diversity of the distributions $\{p(z|\theta)\}$ obtained by sampling the model parameters as $\theta\sim p(\theta|\mathcal{D})$. In fact, the mutual information in (5) equals zero if the distribution $p(z|\theta)$ does not depend on $\theta$, and thus the members of any ensemble $\{p(z|\theta_{k})\}_{k=1}^{K}$ are minimally diverse, all agreeing with each other. Conversely, the measure (5) is large if the distributions $p(z|\theta)$ tend to disagree with each other for different samples $\theta\sim p(\theta|\mathcal{D})$, yielding more diverse ensembles.
Information-theoretically, as indicated by the notation in (5), the difference between the two entropies corresponds to the mutual information $\mathrm{I}(\theta;z|\mathcal{D})$ between the model parameter and the observation $z$ conditioned on a fixed dataset $\mathcal{D}$ (see Fig. 2). This is a measure of the amount of information that we can obtain about the true model parameter $\theta$ by observing another data point $z$, given that we already have access to the dataset $\mathcal{D}$.
We finally note that the mutual information in (5) is computed under the assumed prior $p(\theta)$ and likelihood $p(z|\theta)$. Thus, in order for the conditional mutual information in (5) to be a meaningful measure of epistemic uncertainty, one must assume prior and likelihood to be well-specified.
@@ II-B Generalized Bayesian Learning
Generalized Bayesian learning can be further extended so as to control the reliance of the generalized posterior distribution on the prior distribution. This is done by observing that the generalized posterior $q(\theta|\mathcal{D})$ in (7) is the minimizer of the following information-theoretic objective, known as free energy [11, 6]
@@
where $\text{KL}\left(\cdot\|\cdot\right)$ represents the KL divergence (see Table I). The free energy objective (8) reveals that the generalized posterior distribution arises as a trade-off between fitting the training data, as required by the minimization of the first term in (8), and staying close to the prior, as enforced by the second term in (8) (see Fig. 2).
Replacing the KL divergence with other information-theoretic divergences, such as the Rényi and Tsallis relative entropies [8], yields a generalized form of the free energy objective (8). The minimizer of this objective represents an extension of the notion of generalized posterior, which has been found to be useful in settings with a misspecified prior [11, 23].
@@ II-C Martingale Posterior
over the combined dataset $\mathcal{D}\cup\mathcal{D}^{\prime}$ encompassing both the training set $\mathcal{D}$ and the fictitious data $\mathcal{D}^{\prime}$. The parameters $\theta^{\text{MP}}$ in (10) are random variables due to the stochasticity of the unseen data $\mathcal{D}^{\prime}\sim p(\mathcal{D}^{\prime}|\mathcal{D})$. Samples $\theta^{\text{MP}}$ obtained by drawing independent realizations of unseen data $\mathcal{D}^{\prime}$ are treated as draws from the underlying implicit martingale posterior.
@@ III Information-Theoretic Generalization Bounds
As illustrated in Fig. 1(b), generalization bounds quantify the gap between population loss – the metric ideally minimized during the training of an AI model – and the training loss – the metric actually minimized based on the available data. These bounds provide insights into the training data requirements, potentially offering guidance for algorithm design and data collection strategies. This section briefly reviews generalization bounds that leverage information-theoretic metrics. It will be shown that such bounds provide a theoretical justification for the use of generalized Bayesian learning.
@@ III-A Bounding the Generalization Error
Given a model class, e.g., a class of neural networks, consider a learning algorithm that produces model parameters $\theta$ from the training set $\mathcal{D}=\{z_{1},\ldots,z_{n}\}$, where each data point $z_{i}=(x_{i},y_{i})$ is drawn i.i.d. from a given distribution $p(z)$. The trained model parameters $\theta$ are typically random functions of the training dataset $\mathcal{D}$. In fact, the output of standard training algorithms such as stochastic gradient descent depends on factors such as random initialization and random mini-batch selection. Furthermore, the generalized Bayesian strategies reviewed in the previous section are often implemented to produce samples $\theta\sim q(\theta|\mathcal{D})$ from the generalized posterior, and the martingale posterior, also discussed in Sec. II, generates samples $\theta$ via (10).
@@
To define the generalization error of a training algorithm, with reference to Fig. 1(b), we define the population loss, also known as risk, as the expected value
@@
Intuitively, as illustrated in Fig. 1(b), the generalization error for model $\theta\sim q(\theta|\mathcal{D})$ should decrease as the dataset $\mathcal{D}$ grows in size, since the training loss becomes a more accurate estimate of the population loss. In particular, generalization is known to be problematic when the trained model parameters $\theta$ exhibit an excessive dependence on the specific realization of the training dataset $\mathcal{D}$ – a phenomenon known as overfitting.
@@
where $\mathrm{I}(\mathcal{D};\theta)$ represents the mutual information of the random variables $\mathcal{D},\theta\sim p(\mathcal{D})q(\theta|\mathcal{D})$. Note that the expected value in (14) is also evaluated over the joint distribution of the training data $\mathcal{D}$ and of the output of the training algorithm $\theta\sim q(\theta|\mathcal{D})$. Accordingly, the mutual information $\mathrm{I}(\mathcal{D};\theta)$ between the training set and the learned hypothesis quantifies how much information about the random training data is captured by the algorithm’s output (see Fig. 2). Lower mutual information indicates the algorithm produces more stable outputs across different training sets, suggesting better generalization. Extensions of this bound are reviewed in [9].
For specific algorithm classes, such as stochastic gradient descent with appropriate step sizes, the mutual information $\mathrm{I}(\mathcal{D};\theta)$ grows sub-linearly with $n$, e.g., as $O(\log n)$ [9]. This ensures the right-hand side of the bound (14) vanishes as $n\to\infty$, providing theoretical justification for the effectiveness of the corresponding training algorithms in practice. Moreover, the information-theoretic bound (14) may also guides algorithm design: techniques like gradient noise injection, dropout, and data augmentation can be understood as mechanisms to reduce the mutual information $\mathrm{I}(\mathcal{D};\theta)$, thereby reducing the risk of overfitting and hence the generalization error.
It is also possible to derive high-probability upper bounds on the population loss as a function of the training loss, offering another way to bound the generalization error. Specifically, the probably approximately correct (PAC)-Bayes framework yields bounds of the following form. Fix any constant $\lambda>0$; assume that the loss takes values in the interval $[0,C]$; and choose any prior distribution $p(\theta)$ on the model parameter space, as long as it is independent of the training data. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the random draw of the training set $\mathcal{D}$ of size $n$, the following bound holds [31, 32]

### v2

# Uncertainty Quantification and Data Efficiency in AI: An Information-Theoretic Perspective

@@ I Introduction
  • Quantifying epistemic and predictive uncertainty: In the presence of limited training data, epistemic uncertainty fundamentally limits the precision of any AI model. Uncertainty quantification is thus an essential step to ensure the reliable deployment of AI models in domains such as robotics, telecommunications, and healthcare. We start by reviewing Bayesian learning as a formal framework to quantify epistemic uncertainty in the parameter space, as sketched in Figure 1(a) (with $\theta$ representing the model parameters). We also cover recent “post-Bayesian” advances including generalized Bayesian learning [44] and martingale posteriors [44, 66]. We then review information-theoretic generalization bounds [34]. As illustrated in Figure 1(b), generalization bounds offer insights into the scaling of the average error caused by limitations in the training data. As it will be shown, information-theoretic generalization bounds provide a theoretical justification for the use of generalized Bayesian learning. However, these bounds do not provide any actionable quantification of uncertainty for individual inputs. As exemplified in Figure 1(c), this type of information can be instead obtained via prediction sets calibrated through conformal prediction, which offer distribution-free, finite-sample coverage guarantees [71, 4].
  • Leveraging synthetic data: While uncertainty quantification is a required step whenever training data are in limited supply, in practice one may wish to find ways to augment the dataset with additional data. When real data cannot be collected, or is too expensive to obtain, the only remaining option is to leverage synthetic data. This is potentially feasible by leveraging simulators, or digital twins, of real environments [15, 57, 62] or by adopting powerful general-purpose models such as LLMs [55]. As illustrated in Figure 1(d), synthetic data can be used both to improve training and to enhance the calibration of a pre-trained model. In this context, we review prediction-powered inference, which combines small labeled samples with abundant synthetic predictions for training [5, 83], as well as synthetic-powered predictive inference, which integrates synthetic data for calibration [13, 12].
@@
The definition of the information-theoretic measures used in this article is provided in Table I, while Table II reports a table of acronyms. A summary of information-theoretic relationships surveyed in this paper can be found in Figure 2.
@@ II Generalized Bayesian Learning
Classical Bayesian inference provides a principled framework for incorporating prior knowledge and quantifying epistemic uncertainty through posterior distributions over parameters affecting the data distribution. The principle can be directly applied to the training of AI models, in which a posterior distribution over the model parameters (denoted as $\theta$) is inferred by combining prior information with a likelihood function [69]. As shown in Figure 1(a), this leads to a distribution in the model parameter space, whose extent and shape capture the uncertainty arising from the availability of limited data. In practice, the posterior distribution is approximated using sampling, variational methods, or Laplace approximations [3, 66].
@@
ERM Empirical risk minimization
MCMC Markov chain Monte Carlo
PAC Probably approximately correct
PFN Prior-fitted network
@@
RLHF Reinforcement learning from human feedback
@@
The validity of the posterior distribution hinges on the adherence of prior and likelihood to the ground-truth data distribution. However, prior distributions are often chosen for their analytical tractability, and typical likelihood models fail to capture aspects such as outliers [44, 81]. To obviate these issues, so-called “post-Bayes” methods have been introduced. These move beyond conventional Bayesian inference, while retaining the key goal of quantifying epistemic uncertainty. After reviewing conventional Bayesian learning, this section provides a brief introduction to generalized Bayesian learning, martingale posteriors, and prior-fitted networks as important representative post-Bayes frameworks.
@@ II-A Bayesian Learning
Bayesian inference of parameters $\theta$ begins by specifying a prior distribution $p(\theta)$ and a likelihood function $p(z|\theta)$, where $z$ denotes a data point. The prior $p(\theta)$ should ideally encode available knowledge about the structure and typical values of parameters $\theta$ for which the likelihood $p(z|\theta)$ closely approximates the true data-generating distribution. In supervised learning, the likelihood $p(z|\theta)$ takes the form of a conditional distribution $p(y|x,\theta)$, where $x$ is the input and $y$ is the output variable. More broadly, the likelihood $p(z|\theta)$ defines a model class that encodes the learner’s inductive bias [63]. As we discuss further in Section II-B, it can also be derived from the loss function of a machine learning algorithm.
Given a training dataset (i.e., a multiset with possible multiplicities) $\mathcal{D}=\{z_{1},\ldots,z_{n}\}$, the posterior distribution is given by
@@
Unlike conventional frequentist learning, which returns a single model parameter $\theta$, the posterior distribution (1) directly quantifies epistemic uncertainty through the spread of probability mass in the parameter space. Wide posterior distributions indicate high uncertainty about which parameters best explain the data, signaling insufficient training samples. As the training dataset grows in size, i.e., in the limit $n\to\infty$, consistency results show that, under appropriate conditions, the posterior concentrates around the optimal parameter $\theta^{*}$, as illustrated in Figure 1(a). Specifically, the optimal parameter $\theta^{*}$ provides the best fit for the true data distribution in terms of the Kullback-Leibler (KL) divergence (see Table I) [50].
Trained predictive or generative models can be obtained by drawing one or more samples from the distribution $p(\theta|\mathcal{D})$, typically using Markov chain Monte Carlo (MCMC) methods [3]. For example, the unadjusted Langevin algorithm produces the sequence of samples
@@
The quantity (5) is a measure of the diversity of the distributions $\{p(z|\theta)\}$ obtained by sampling the model parameters as $\theta\sim p(\theta|\mathcal{D})$. In fact, the mutual information in (5) equals zero if the distribution $p(z|\theta)$ does not depend on $\theta$, and thus the members of any ensemble $\{p(z|\theta_{k})\}_{k=1}^{K}$ are minimally diverse, all agreeing with each other. Conversely, the measure (5) is large if the distributions $p(z|\theta)$ tend to disagree with each other for different samples $\theta\sim p(\theta|\mathcal{D})$, yielding more diverse ensembles. Note that the mutual information in (5) is computed under the assumed prior $p(\theta)$ and likelihood $p(z|\theta)$. Thus, in order for the conditional mutual information in (5) to be a meaningful measure of epistemic uncertainty, one must assume prior and likelihood to be well-specified.
Information-theoretically, as indicated by the notation in (5), the difference between the two entropies corresponds to the mutual information $\mathrm{I}(\theta;z|\mathcal{D})$ between the model parameter and the observation $z$ conditioned on a fixed dataset $\mathcal{D}$ (see Figure 2). This is a measure of the amount of information that we can obtain about the true model parameter $\theta$ by observing another data point $z$, given that we already have access to the dataset $\mathcal{D}$.
This epistemic uncertainty measure, given by the mutual information (5) between parameters and data, plays a central role in Bayesian active learning, where it is used to select the most informative data points for labeling [37]. Beyond active learning, mutual information-based epistemic uncertainty estimates have been leveraged, among other applications, for out-of-distribution detection [48] and safe decision-making in reinforcement learning [22].
@@ II-B Generalized Bayesian Learning
Generalized Bayesian learning can be further extended so as to control the reliance of the generalized posterior distribution on the prior distribution. This is done by observing that the generalized posterior $q(\theta|\mathcal{D})$ in (7) is the minimizer of the following information-theoretic objective, known as free energy [44, 66, 56]
@@
where $\text{KL}\left(\cdot\|\cdot\right)$ represents the KL divergence (see Table I). The free energy objective (8) reveals that the generalized posterior distribution arises as a trade-off between fitting the training data, as required by the minimization of the first term in (8), and staying close to the prior, as enforced by the second term in (8) (see Figure 2).
Replacing the KL divergence with other information-theoretic divergences, such as the Rényi and Tsallis relative entropies [66], yields a generalized form of the free energy objective (8). The minimizer of this objective represents an extension of the notion of generalized posterior, which has been found to be useful in settings with a misspecified prior [44, 81, 21].
The free energy (8) provides a principled framework for understanding regularized objectives in LLM alignment. Notably, in the context of reinforcement learning from human feedback (RLHF), we interpret $q$ as the LLM being optimized being optimized; $p$ as a reference policy corresponding to the pre-trained LLM; and the training loss $L_{\text{train}}$ as the negative reward assigned by a learned reward model. The KL term thus penalizes deviation from the reference distribution, preventing the policy from exploiting spurious correlations in the imperfect reward signal, while preserving the general capabilities acquired during pre-training. The trade-off between maximizing expected reward (the first term) and remaining close to a trusted prior policy (the second term) is dictated by the parameter $\lambda$, whose selection is central to RLHF-based alignment methods [53].
@@ II-C Martingale Posterior
over the combined dataset $\mathcal{D}\cup\mathcal{D}^{\prime}$ encompassing both the training dataset $\mathcal{D}$ and the fictitious data $\mathcal{D}^{\prime}$. The parameters $\theta^{\text{MP}}$ in (10) are random variables due to the stochasticity of the unseen data $\mathcal{D}^{\prime}\sim p(\mathcal{D}^{\prime}|\mathcal{D})$. Samples $\theta^{\text{MP}}$ obtained by drawing independent realizations of unseen data $\mathcal{D}^{\prime}$ are treated as draws from the underlying implicit martingale posterior.
@@
II-D Prior-Fitted Networks
As discussed in this section, a central challenge in deploying Bayesian learning in practice is the computational cost of posterior inference. Standard methods, such as MCMC sampling methods or variational inference [3, 66], require iterative optimization or sampling procedures, such as (2), that must be repeated from scratch for every new dataset. This difficulty is compounded when the prior encodes complex, domain-specific generative assumptions, e.g., priors defined through structural causal models or simulators.
Prior-fitted networks (PFNs) [52] recasts Bayesian inference as a supervised meta-learning problem, leveraging the in-context learning (ICL) capabilities of modern sequence models. Specifically, a transformer network is trained on a large collection of synthetic datasets, each drawn from the given prior. Given a training dataset of input-output pairs as context, the network learns to predict the posterior predictive distribution for held-out query points in a single forward pass without any iterative adaptation at test time. This way, the cost of learning is shifted from test time to a one-time pretraining phase, amortizing the cost across an arbitrary number of Bayesian inference tasks.
Applications of PFNs include Bayesian optimization [51], tabular models [35], and causal inference [10]. Further discussions on the use of synthetic data can be found in Section V.
@@ III Information-Theoretic Generalization Bounds
As illustrated in Figure 1(b), generalization bounds quantify the gap between population loss – the metric ideally minimized during the training of an AI model – and the training loss – the metric actually minimized based on the available data. These bounds provide insights into the training data requirements, potentially offering guidance for algorithm design and data collection strategies. This section briefly reviews generalization bounds that leverage information-theoretic metrics. It will be shown that such bounds provide a theoretical justification for the use of generalized Bayesian learning.
@@ III-A Bounding the Generalization Error
Given a model class, e.g., a class of neural networks, consider a learning algorithm that produces model parameters $\theta$ from the training dataset $\mathcal{D}=\{z_{1},\ldots,z_{n}\}$, where each data point $z_{i}=(x_{i},y_{i})$ is drawn i.i.d. from a given distribution $p(z)$. The trained model parameters $\theta$ are typically random functions of the training dataset $\mathcal{D}$. In fact, the output of standard training algorithms such as stochastic gradient descent depends on factors such as random initialization and random mini-batch selection. Furthermore, the generalized Bayesian strategies reviewed in the previous section are often implemented to produce samples $\theta\sim q(\theta|\mathcal{D})$ from the generalized posterior, and the martingale posterior, also discussed in Section II, generates samples $\theta$ via (10).
@@
To define the generalization error of a training algorithm, with reference to Figure 1(b), we define the population loss, also known as risk, as the expected value
@@
Intuitively, as illustrated in Figure 1(b), the generalization error for model $\theta\sim q(\theta|\mathcal{D})$ should decrease as the dataset $\mathcal{D}$ grows in size, since the training loss becomes a more accurate estimate of the population loss. In particular, generalization is known to be problematic when the trained model parameters $\theta$ exhibit an excessive dependence on the specific realization of the training dataset $\mathcal{D}$ – a phenomenon known as overfitting.
III-B Bounding the Average Generalization Error
@@
where $\mathrm{I}(\mathcal{D};\theta)$ represents the mutual information of the random variables $\mathcal{D},\theta\sim p(\mathcal{D})q(\theta|\mathcal{D})$. Note that the expected value in (14) is also evaluated over the joint distribution of the training data $\mathcal{D}$ and of the output of the training algorithm $\theta\sim q(\theta|\mathcal{D})$. Accordingly, the mutual information $\mathrm{I}(\mathcal{D};\theta)$ between the training dataset and the learned hypothesis quantifies how much information about the random training data is captured by the algorithm’s output (see Figure 2). Lower mutual information indicates the algorithm produces more stable outputs across different training sets, suggesting better generalization. Extensions of this bound are reviewed in [34].
For specific algorithm classes, such as stochastic gradient descent with appropriate step sizes, the mutual information $\mathrm{I}(\mathcal{D};\theta)$ grows sub-linearly with $n$, e.g., as $O(\log n)$ [34]. This ensures the right-hand side of the bound (14) vanishes as $n\to\infty$, providing theoretical justification for the effectiveness of the corresponding training algorithms in practice. Moreover, the information-theoretic bound (14) may also guide algorithm design: techniques like gradient noise injection, dropout, and data augmentation can be understood as mechanisms to reduce the mutual information $\mathrm{I}(\mathcal{D};\theta)$, thereby reducing the risk of overfitting and hence the generalization error. Finally, exact expressions for the average generalization error have been derived for the Gibbs posterior (7) [2].
III-C Bounding the Generalization Error with High Probability
It is also possible to derive high-probability upper bounds on the population loss as a function of the training loss, offering another way to bound the generalization error. Specifically, the probably approximately correct (PAC)-Bayes framework yields bounds of the following form. Fix any constant $\lambda>0$; assume that the loss takes values in the interval $[0,C]$; and choose any prior distribution $p(\theta)$ on the model parameter space, as long as it is independent of the training data. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the random draw of the training dataset $\mathcal{D}$ of size $n$, the following bound holds [18, 1]